10x-10y=0
5x-4y=8


Answer by schmidtmomto6(24)

10x-10y=0
5x-4y=8
Multiply the bottom equation by -2: -10x+8y=-16.
Now add this second equation to the first one:
10x-10y = 0
-10x+ 8y =-16
---------------
0-2y = -16, so y=8
Stick this value for y back in one of the original equations to solve for x:
10x-10(8)=0, 10x-80=0, 10x=80, x=8
Hope this helps! =)